Deduce the dimensional formula for modulus of elasticity(Y).  

(a) ML-1T-2

(b) ML-3T-2  

(c) ML-2T-2  

(d) ML-1T-3



1
Expert's answer
2021-02-10T10:06:45-0500
"Y=\dfrac{FL_0}{A\Delta L}=\dfrac{\dfrac{L}{T^2}ML}{L^3}=ML^{-1}T^{-2}."

Answer:

(a) "ML^{-1}T^{-2}"